1. Define PV, FV, PVA, FVA, and PV . using an example with close relevance to you. The example may be professionally (e.g., issuing bonds and stocks of your corporation) or personally (e.g., investing your retirement assets) related to you.
Answer #1
1) PV:
Refers to the Present Value of a Future
Lump sum to be received when discounted
at the appropriate rate.
The formula for finding PV of a lump sum is
PV = FV/(1+r)^n
where, r = the discount rate and n = the # of
years or periods
Example:
I have an obligation to make a payment of
$5000 to my friend after 5 years. If I wish to
settle the amount today I need pay only its
PV when discounted at a rate agreeable to
my friend.
If such rate is 5%, the amount to be paid
today = 5000/1.05^ 5 = $         3,918
2) FV:
FV is the opposite of PV and represents the
future value of a lump sum today.
FV = PV*(1+r)^n
Example:
I have made a deposit of $5000 in a bank
which pays interest at 5%, compounded
annually. If the maturity is 5 years, the
value of the deposit with interest will be
5000*1.05^5 = $         6,381
3) PVA:
PVA means PV of annuity. Annuity is series
of equal payments for a specified number
of periods.
The PV of annuity = [(1+r)^n-1]/[r*(1+r)^n]
I needed an annuity of $500 (yearly) for 25 years
for which I approached an Insurance co.
The company asked me to pay the PV of the
annuity, the discount rate being 3% p.a.
The PV of annuity or the amount payable today to purchase the annuity = 500*(1.03^25-1)/(0.03*1.03^25) = $         8,707
4) FVA is the opposite of PVA.
It is given by the formula:
FVA = [(1+r)^n-1]/r
Example:
I plan to save $500 for 5 years at the end of each
year. If I can earn an interst rate of 3% compounded
annually, the FV of this annuity will be:
500*(1.03^5-1)/0.03 = $         2,655